# Stereo-GS: Multi-View Stereo Vision Model for Generalizable 3D Gaussian Splatting Reconstruction

### v1

@@ 4.2. Disentangled 3D Gaussians prediction
Existing methods (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024) for generalizable 3D Gaussian prediction typically take multi-view images as input and entangle the prediction of 3D Gaussian geometry with other appearance parameters. Nonetheless, the absence of geometry prior knowledge leads to predicting the 3D Gaussian positions and learning multi-view correspondence reasoning in an implicit manner. Such implicit predictions can cause the regression to solely rely on data-driven priors for learning multi-view correspondence reasoning and geometry estimation, which is time-consuming and can result in unreliable 3D Gaussian predictions.
We propose a disentangled two-stage training approach to predict 3D Gaussians from multi-view images, where the geometry and appearance of the 3D Gaussians are learned separately. Unlike previous methods (Szymanowicz et al., 2024; Zou et al., 2024; Tang et al., 2024a), which are limited to fixed input configurations, our method is flexible and can handle arbitrary multi-view images, significantly broadening its applicability.
@@
A key advantage of our Stereo-GS is its pose-free nature, which ensures generalizability and robustness for direct 3D reconstruction tasks. In contrast, existing methods (Tang et al., 2024a) typically assume the availability of precise camera parameters during testing. This reliance on camera parameters not only limits their applicability to real-world scenarios but also makes them highly sensitive to inaccuracies in the provided parameters. For example, imperfect or noisy camera poses, especially in synthetic multi-view image settings, can lead to suboptimal reconstructions and degraded results.
@@
In the second stage, we train our model with the Gaussian prediction head for 3DGS appearance prediction. Each pixel in the output feature map represents a 3D Gaussian appearance attribute, similar to SplatterImage (Szymanowicz et al., 2024). The output feature map consists of $11$ channels, encoding the appearance information of 3D Gaussians, including scaling $s$, rotation $q$, opacity $\sigma$ and RGB color $c$. We concatenate these appearance attributes with the previously predicted point maps $P$, which represent the 3D Gaussian positions $\mu$, to construct GS-maps that align with the standard 3D Gaussian parameters.
For the output GS-maps, we clamp the point-maps position $P$ into $[-1.0,1.0]^{3}$. Color $c$ and opacity $\alpha$ are activated with the sigmoid function. Rotation $r$ is $\ell_{2}$ normalized. Scaling $s$ is activated with the softplus function and multiplied with a small coefficient $0.1$. We transform each GS-map into a set of Gaussians and directly merge them as the final 3D Gaussians, which are used to render images as novel views during training.
@@
After the two-stage training of geometry and appearance, our Stereo-GS model can directly generate multi-view GS-maps, efficiently representing the 3DGS objects. To further enhance the reconstruction quality, we introduce a U-Net-based refinement network to refine the GS-maps. We follow the way in MASt3R (Leroy et al., 2024) to make the local stereo model to predict matching features intrinsically contain rich geometric information from the pair-wise images. Such geometric information can serve as a prior for multi-view correspondence learning during the refinement network training.
In specific, we directly concatenate the multi-view GS-maps with the matching features and RGB images as the input to the refinement network. The refinement network is a U-Net network and constructed with residual layers (He et al., 2016) and self-attention layers (Vaswani, 2017) following previous works (Ho et al., 2020; Metzer et al., 2023; Szymanowicz et al., 2024). Self-attention layers are incorporated at the deeper layers where the input multi-view feature embedding resolution is down-sampled. We flatten and concatenate the multi-view features before applying self-attention layers, facilitating information propagation across multiple views, similar to prior multi-view diffusion models (Wang and Shi, 2023; Shi et al., 2023b). The output of the refinement network is added to the input Gaussian maps as a residual connection.
@@ 4.3. Training
Our training contains two stages to optimize the predicted 3DGS geometry and appearance. The training is conducted on the large-scale 3D object (Deitke et al., 2023) dataset for generalizability and robustness. In the first stage, we focus on training the Stereo-GS model with the point prediction head for geometry prediction. To ensure accurate ground truth geometry for training, we randomly sample 3D points from the surface of the 3D object to construct a ground truth point cloud $\hat{S}$. This approach provides a reliable reference for learning precise multi-view point-maps. During training, we employ a regression-based method to optimize the prediction of multi-view point-maps. Specifically, we randomly sample points from the predicted point-maps to form a predicted point cloud $S$. To measure the alignment between the predicted and ground truth point clouds, we use the Chamfer Distance loss:
@@
(6) $\mathcal{L}_{\text{depth}}=\alpha\cdot|D-\hat{D}|+$
$\beta\cdot(|\partial_{x}D-\partial_{x}\hat{D}|+|\partial_{y}D-\partial_{y}\hat{D}|),$
where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions. Following UniMatch (Xu et al., 2023), we use $\alpha=20$ and $\beta=20$.
In the second stage, we train our Stereo-GS model with the Gaussian prediction head and refinement network for appearance prediction and novel view synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS (Zhang et al., 2018) losses between rendered images and ground truth:
@@ 4.4. Implementation
We train our model on 8 NVIDIA V100 (32G) GPUs for about 2 days. Previous methods (Tang et al., 2024a; Hong et al., 2023; Xu et al., 2024b) which requires higher-end A100 (80G) GPUs for more than 3000 hours. Our method only needs lower-end V100 GPUs for slightly above 300 hours, which significantly reduces the training hours to 1/10 even with lower-end GPUs. Each GPU uses a batch size of 6 with bfloat16 precision, resulting in a total batch size of 48. We set 4 images as the input images by default and we randomly select 8 camera views in each batch. We use the first 4 views as input and all 8 views for supervision. We render the output GS-maps at $512\times 512$ resolution for mean square error loss and resized to $256\times 256$ for LPIPS loss. The AdamW (Loshchilov, 2017) optimizer is adopted with the learning rate of $4\times 10^{-4}$, weight decay of $0.05$, and betas of $(0.9,0.95)$. The learning rate is set to 00 for a cosine restart annealing learning rate scheduler every 20 epochs and the training can be finished in 200 epochs.
Figure 3. Multi-view reconstruction. Given the same multi-view inputs, the standard GS (Kerbl et al., 2023) totally fails to render novel view images. SplatterImage hardly reconstructs the 3D objects under the multiple view images. Although LGM (Tang et al., 2024a) can reconstruct finer geometric structures and appearance details, it still faces challenges in maintaining consistency and avoiding artifacts. Our method can generate both high quality geometry and appearance for the 3D objects.
@@ 5. Experiments
Our training dataset consists of multi-view images rendered from the Objaverse (Deitke et al., 2023) dataset. Unlike existing methods that typically require a large-scale dataset of the full set of Objaverse with around 80K objects (Hong et al., 2023), or require around one million rendering views (Tang et al., 2024a). We utilize the LVIS annotation to obtain a subset of Objaverse (Deitke et al., 2023), which contains around $45$k high-quality objects. For each object in the dataset, we render 512 × 512 images and depth maps from 32 random viewpoints. Our datasets contain around 144k rendering views in total, which is sufficient for our model training to obtain superior performance and reduce the reliance on large-scale datasets. We evaluate the qualitative and quantitative performance using Google Scanned Objects (GSO) (Downs et al., 2022) datasets. The GSO dataset comprises around 1,000 objects, from which we randomly select 100 objects for our evaluation set. Specifically, we render 16 images of each object in the GSO evaluation set in an orbiting trajectory with uniform azimuths varying positive elevations in $\{0^{\circ},20^{\circ}\}$ for sampling on the top semi-sphere of an object.
@@ 5.1. Results on multi-view reconstruction
We compare our method with Gaussian Splatting (Kerbl et al., 2023), SplatterImage (Szymanowicz et al., 2024) and LGM (Tang et al., 2024a) on the GSO dataset. For all baselines, we use four input views with camera elevation angle = 0 and azimuths degrees = $\{0,90,180,270\}$ to cover the entire object and evaluate the reconstruction quality on the remaining 12 views. SplatterImage (Szymanowicz et al., 2024) was originally designed to handle multi-view reconstruction using only two input views. When all four input view images are used, it results in corrupted reconstructions. Therefore, to align with the original implementation, we only use the first two views from the input images.
We use PSNR, SSIM, and LPIPS (Zhang et al., 2018) to measure the appearance reconstruction quality. Following previous methods (Tang and Tan, 2018; Xu et al., 2023), we use Absolute Relative difference (Abs Rel), Squared Relative difference (Sq Rel), and Root Mean Squared Error (RMSE) for evaluation of the rendered depth maps to measure geometry reconstruction quality. All comparisons are conducted at a resolution of $256$. As shown in Figure 3, our reconstructions accurately represent the geometric structures without any visible floaters and capture appearance details with higher quality than the baseline methods. Table 1 and Table 2 demonstrate the quantitative reconstruction results, our method outperforms all baselines across both appearance and geometry metrics with obvious margins. Though our method takes a slightly higher inference speed than LGM (Tang et al., 2024a), it shows a balance between high reconstruction quality and acceptable inference efficiency. We provide more results for real-world objects and scene scenarios in the supplementary material to further demonstrate the scalability of our method.
Table 1. Appearance reconstruction quality evaluation for multi-view reconstruction on Google Scanned Objects datasets (Downs et al., 2022).
Methods PSNR↑ SSIM↑ LPIPS↓ Time GS (Kerbl et al., 2023) 17.81 0.6946 0.2948 270 sec SplatterImage (Szymanowicz et al., 2024) 18.05 0.8221 0.1764 0.32 sec LGM (Tang et al., 2024a) 23.79 0.8791 0.1035 1.67 sec Ours 27.12 0.9352 0.0534 2.62 sec
Table 2. Geometry reconstruction quality evaluation for multi-view reconstruction on Google Scanned Objects datasets (Downs et al., 2022).
Method Abs Rel↓ Sq Rel↓ RMSE↓ GS (Kerbl et al., 2023) 0.4570 0.1982 0.5031 SplatterImage (Szymanowicz et al., 2024) 0.3223 0.1178 0.3779 LGM (Tang et al., 2024a) 0.2695 0.0758 0.1973 Ours 0.1112 0.0238 0.1175
@@
Figure 5. Single Image-to-3D generation with 4 input views and 6 input views. We use V3D (Chen et al., 2024a) to generate multi-view images, with the input image scaled up for better visualization and positioned in the upper-left corner of each input views group.
@@ 5.2. Results on single Image-to-3D generation
We utilize the recently proposed V3D (Chen et al., 2024a) as our multi-view diffusion model since its superior performance in generating high-quality multi-view consistent images as the multi-view input for our model. The camera elevation is fixed to 0, and azimuths are set to $[0,80,180,280]$ degree for the four input views condition and $[0,60,120,180,240,300]$ degree for the six input views condition. The single input image is preprocessed for background removal using the $\operatorname{U^{2}Net}$ (Qin et al., 2020) method.
We compare our method with four recent baselines (Szymanowicz et al., 2024; Tang et al., 2024b; Zou et al., 2024; Tang et al., 2024a) that are capable of generating 3D Gaussians from a single image input. As shown in Figure 4, we display the image rendered by different methods for comparison. SplatterImage (Szymanowicz et al., 2024) shows poor novel view synthesis results. DreamGaussian (Tang et al., 2024b) is an optimization-based method and shows various geometry artifacts. TriplaneGaussian (Zou et al., 2024) learns to generate flat geometry, leading to unrealistic results and missing content details. LGM (Tang et al., 2024a) shows multi-view inconsistency, resulting in geometry artifacts and blurry texture. Our method generates consistent results with higher reconstruction quality in both appearance and geometry. We also quantitatively evaluate the reconstruction result on the GSO data by using PSNR, SSIM, and LPIPS (Zhang et al., 2018) metrics as shown in Table 3.
Table 3. Single image-to-3D generation on GSO dataset. Combined with an image-to-multiview diffusion model (Shi et al., 2023a), our method can be used for single image-to-3D generation.
Method PSNR↑ SSIM↑ LPIPS↓ DreamGaussian (Tang et al., 2024b) 18.23 0.8161 0.1897 Splatter-Image (Szymanowicz et al., 2024) 16.06 0.7804 0.3289 TriplaneGaussian (Zou et al., 2024) 16.48 0.8085 0.2567 LGM (Tang et al., 2024a) 17.04 0.8173 0.2274 Ours 19.31 0.8238 0.1412
@@ 5.3. Ablation study
Besides the 4 input image settings, we also train our model with 6 input images and generate 3D Gaussians under the supervision of 8 images. As illustrated in Figure 5, our method achieves faithful 3DGS reconstruction using both 4 and 6 input images. Notably, qualitative results with 6 input images demonstrate greater consistency, indicating that leveraging additional images enhances reconstruction quality. Previous methods typically use a fixed number of input images (Szymanowicz et al., 2024; Tang et al., 2024a). In contrast, our model can reconstruct 3D objects with a larger number of input images, allowing more synthetic images as input to enhance the overall reconstruction quality. We provide more details about the qualitative and quantitative results in the supplementary materials.
@@
We compare the training efficiency with LGM (Tang et al., 2024a). Our experiments are conducted using our default training resources, which include 8 NVIDIA V100 (32GB) GPUs, and we train on the Objaverse (Deitke et al., 2023) LVIS subset. To make a fair comparison, we set the input to 4 images and set the output Gaussians resolution to $256\times 256=65536$ for both methods. Specifically, we add one more upsampling layer for LGM to match the input resolution. As shown in Figure 6 (a), our method converges within 200 epochs. In contrast, LGM (Tang et al., 2024a) still requires significantly more training time to reach convergence.
Additionally, we experimented with a smaller output Gaussian resolution of $128\times 128=16384$, which follows the original asymmetric design of LGM (Tang et al., 2024a). The smaller size of predicted 3D Gaussians enables a more effective batch size, allowing our method to converge faster within around 100 epochs. Meanwhile, LGM(Tang et al., 2024a) still requires more than 200 epochs to complete training. Both experiments indicate that our method successfully utilizes the disentangled framework to accelerate the reconstruction model regression.

### v2

@@ 4.2. Disentangled 3D Gaussians prediction
Existing methods (tang2024lgm; xu2024grm; gslrm2024) for generalizable 3D Gaussian prediction typically take multi-view images as input and entangle the prediction of 3D Gaussian geometry with other appearance parameters. Nonetheless, the absence of geometry prior knowledge leads to predicting the 3D Gaussian positions and learning multi-view correspondence reasoning in an implicit manner. Such implicit predictions can cause the regression to solely rely on data-driven priors for learning multi-view correspondence reasoning and geometry estimation, which is time-consuming and can result in unreliable 3D Gaussian predictions.
We propose a disentangled two-stage training approach to predict 3D Gaussians from multi-view images, where the geometry and appearance of the 3D Gaussians are learned separately. Unlike previous methods (szymanowicz2024splatter; zou2024triplane; tang2024lgm), which are limited to fixed input configurations, our method is flexible and can handle arbitrary multi-view images, significantly broadening its applicability.
@@
A key advantage of our Stereo-GS is its pose-free nature, which ensures generalizability and robustness for direct 3D reconstruction tasks. In contrast, existing methods (tang2024lgm) typically assume the availability of precise camera parameters during testing. This reliance on camera parameters not only limits their applicability to real-world scenarios but also makes them highly sensitive to inaccuracies in the provided parameters. For example, imperfect or noisy camera poses, especially in synthetic multi-view image settings, can lead to suboptimal reconstructions and degraded results.
@@
In the second stage, we train our model with the Gaussian prediction head for 3DGS appearance prediction. Each pixel in the output feature map represents a 3D Gaussian appearance attribute, similar to SplatterImage (szymanowicz2024splatter). The output feature map consists of $11$ channels, encoding the appearance information of 3D Gaussians, including scaling $s$, rotation $q$, opacity $\sigma$ and RGB color $c$. We concatenate these appearance attributes with the previously predicted point maps $P$, which represent the 3D Gaussian positions $\mu$, to construct GS-maps that align with the standard 3D Gaussian parameters.
For the output GS-maps, we clamp the point-maps position $P$ into $1.0,1.0]^{3}$. Color $c$ and opacity $\alpha$ are activated with the sigmoid function. Rotation $r$ is $\ell_{2}$ normalized. Scaling $s$ is activated with the softplus function and multiplied with a small coefficient $0.1$. We transform each GS-map into a set of Gaussians and directly merge them as the final 3D Gaussians, which are used to render images as novel views during training.
@@
After the two-stage training of geometry and appearance, our Stereo-GS model can directly generate multi-view GS-maps, efficiently representing the 3DGS objects. To further enhance the reconstruction quality, we introduce a U-Net-based refinement network to refine the GS-maps. We follow the way in MASt3R (mast3r_eccv24) to make the local stereo model to predict matching features intrinsically contain rich geometric information from the pair-wise images. Such geometric information can serve as a prior for multi-view correspondence learning during the refinement network training. In addition, camera pose information is often available in synthetic data or multi-view generative models such as Mvdream (shi2023mvdream) or V3D(chen2024v3d), which can also be leveraged by our refinement network to produce more fine-grained results.
In specific, we directly concatenate the multi-view GS-maps with the matching features and RGB images as the input to the refinement network. Additionally, camera poses can be encoded as Plücker ray embeddings (xu2023dmv3d) and optionally concatenated with the multi-view inputs. The refinement network is a U-Net network and constructed with residual layers (he2016deep) and self-attention layers (vaswani2017attention) following previous works (ho2020denoising; metzer2023latent; szymanowicz2024splatter). Self-attention layers are incorporated at the deeper layers where the input multi-view feature embedding resolution is down-sampled. We flatten and concatenate the multi-view features before applying self-attention layers, facilitating information propagation across multiple views, similar to prior multi-view diffusion models (wang2023imagedream; shi2023mvdream). The output of the refinement network is added to the input Gaussian maps as a residual connection.
@@ 4.3. Training
Our training contains two stages to optimize the predicted 3DGS geometry and appearance. The training is conducted on the large-scale 3D object (deitke2023objaverse) dataset for generalizability and robustness. In the first stage, we focus on training the Stereo-GS model with the point prediction head for geometry prediction. To ensure accurate ground truth geometry for training, we randomly sample 3D points from the surface of the 3D object to construct a ground truth point cloud $\hat{S}$. This approach provides a reliable reference for learning precise multi-view point-maps. During training, we employ a regression-based method to optimize the prediction of multi-view point-maps. Specifically, we randomly sample points from the predicted point-maps to form a predicted point cloud $S$. To measure the alignment between the predicted and ground truth point clouds, we use the Chamfer Distance loss:
@@
(6) $D^|+\displaystyle\mathcal{L}_{\text{depth}}=\alpha\cdot$|D-$\hat${D}|+
[…]
where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions. Following UniMatch (xu2023unifying), we use $\alpha=20$ and $\beta=20$.
In the second stage, we train our Stereo-GS model with the Gaussian prediction head and refinement network for appearance prediction and novel view synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS (Zhang_Isola_Efros_Shechtman_Wang_2018) losses between rendered images and ground truth:
@@ 4.4. Implementation
We train our model on 8 NVIDIA V100 (32G) GPUs for about 2 days. Previous methods (tang2024lgm; hong2023lrm; xu2024grm) which requires higher-end A100 (80G) GPUs for more than 3000 hours. Our method only needs lower-end V100 GPUs for slightly above 300 hours, which significantly reduces the training hours to 1/10 even with lower-end GPUs. Each GPU uses a batch size of 6 with bfloat16 precision, resulting in a total batch size of 48. We set 4 images as the input images by default and randomly select 8 camera views in each batch. The AdamW (loshchilov2017decoupled) optimizer is adopted with the learning rate of $4\times 10^{-4}$, weight decay of $0.05$, and betas of $(0.9,0.95)$.
Figure 3. Multi-view reconstruction. Given the same multi-view inputs, the standard GS (kerbl3Dgaussians) totally fails to render novel view images. SplatterImage hardly reconstructs the 3D objects under the multiple view images. Although LGM (tang2024lgm) can reconstruct finer geometric structures and appearance details, it still faces challenges in maintaining consistency and avoiding artifacts. Our method can generate both high quality geometry and appearance for the 3D objects.
@@ 5. Experiments
Our training dataset consists of multi-view images rendered from the Objaverse (deitke2023objaverse) dataset. Unlike existing methods that typically require a large-scale dataset of the full set of Objaverse with around 80K objects (hong2023lrm), or require around one million rendering views (tang2024lgm). We utilize the LVIS annotation to obtain a subset of Objaverse (deitke2023objaverse), which contains around 4545k high-quality objects. For each object in the dataset, we render 512 × 512 images and depth maps from 32 random viewpoints. Our datasets contain around 144k rendering views in total, which is sufficient for our model training to obtain superior performance and reduce the reliance on large-scale datasets. We evaluate the qualitative and quantitative performance using Google Scanned Objects (GSO) (downs2022google) datasets. The GSO dataset comprises around 1,000 objects, from which we randomly select 100 objects for our evaluation set. Specifically, we render 16 images of each object in the GSO evaluation set in an orbiting trajectory with uniform azimuths varying positive elevations in $\{0^{\circ},20^{\circ}\}$ for sampling on the top semi-sphere of an object.
@@ 5.1. Results on multi-view reconstruction
We compare our method with Gaussian Splatting (kerbl3Dgaussians), SplatterImage (szymanowicz2024splatter) and LGM (tang2024lgm) on the GSO dataset. For all baselines, we use four input views with camera elevation angle = 0 and azimuths degrees = $\{0,90,180,270\}$ to cover the entire object and evaluate the reconstruction quality on the remaining 12 views. SplatterImage (szymanowicz2024splatter) was originally designed to handle multi-view reconstruction using only two input views. When all four input view images are used, it results in corrupted reconstructions. Therefore, to align with the original implementation, we only use the first two views from the input images.
We use PSNR, SSIM, and LPIPS (Zhang_Isola_Efros_Shechtman_Wang_2018) to measure the appearance reconstruction quality. Following previous methods (tang2018ba; xu2023unifying), we use Absolute Relative difference (Abs Rel), Squared Relative difference (Sq Rel), and Root Mean Squared Error (RMSE) for evaluation of the rendered depth maps to measure geometry reconstruction quality. All comparisons are conducted at a resolution of $256$. As shown in Figure 3, our reconstructions accurately represent the geometric structures without any visible floaters and capture appearance details with higher quality than the baseline methods. Table 1 and Table 2 demonstrate the quantitative reconstruction results, our method outperforms all baselines across both appearance and geometry metrics with obvious margins. Though our method takes a slightly higher inference speed than LGM (tang2024lgm), it shows a balance between high reconstruction quality and acceptable inference efficiency. We provide more results for real-world objects and scene scenarios in the supplementary material to further demonstrate the scalability of our method.
Table 1. Appearance reconstruction quality evaluation for multi-view reconstruction on Google Scanned Objects datasets (downs2022google).
Methods PSNR↑ SSIM↑ LPIPS↓ Time GS (kerbl3Dgaussians) 17.81 0.6946 0.2948 270 sec SplatterImage (szymanowicz2024splatter) 18.05 0.8221 0.1764 0.32 sec LGM (tang2024lgm) 23.79 0.8791 0.1035 1.67 sec Ours 27.12 0.9352 0.0534 2.62 sec
Table 2. Geometry reconstruction quality evaluation for multi-view reconstruction on Google Scanned Objects datasets (downs2022google).
Method Abs Rel↓ Sq Rel↓ RMSE↓ GS (kerbl3Dgaussians) 0.4570 0.1982 0.5031 SplatterImage (szymanowicz2024splatter) 0.3223 0.1178 0.3779 LGM (tang2024lgm) 0.2695 0.0758 0.1973 Ours 0.1112 0.0238 0.1175
@@
Figure 5. Single Image-to-3D generation with 4 input views and 6 input views. We use V3D (chen2024v3d) to generate multi-view images, with the input image scaled up for better visualization and positioned in the upper-left corner of each input views group.
@@ 5.2. Results on single Image-to-3D generation
We utilize the recently proposed V3D (chen2024v3d) as our multi-view diffusion model since its superior performance in generating high-quality multi-view consistent images as the multi-view input for our model. The camera elevation is fixed to 0, and azimuths are set to $[0,80,180,280]$ degree for the four input views condition and $[0,60,120,180,240,300]$ degree for the six input views condition. The single input image is preprocessed for background removal using the $\operatorname{U^{2}Net}$ (Qin_2020_PR) method.
We compare our method with four recent baselines (szymanowicz2024splatter; tang2023dreamgaussian; zou2024triplane; tang2024lgm) that are capable of generating 3D Gaussians from a single image input. As shown in Figure 4, we display the image rendered by different methods for comparison. SplatterImage (szymanowicz2024splatter) shows poor novel view synthesis results. DreamGaussian (tang2023dreamgaussian) is an optimization-based method and shows various geometry artifacts. TriplaneGaussian (zou2024triplane) learns to generate flat geometry, leading to unrealistic results and missing content details. LGM (tang2024lgm) shows multi-view inconsistency, resulting in geometry artifacts and blurry texture. Our method generates consistent results with higher reconstruction quality in both appearance and geometry. We also quantitatively evaluate the reconstruction result on the GSO data by using PSNR, SSIM, and LPIPS (Zhang_Isola_Efros_Shechtman_Wang_2018) metrics as shown in Table 3.
Table 3. Single image-to-3D generation on GSO dataset. Combined with an image-to-multiview diffusion model (shi2023zero123++), our method can be used for single image-to-3D generation.
Method PSNR↑ SSIM↑ LPIPS↓ DreamGaussian (tang2023dreamgaussian) 18.23 0.8161 0.1897 Splatter-Image (szymanowicz2024splatter) 16.06 0.7804 0.3289 TriplaneGaussian (zou2024triplane) 16.48 0.8085 0.2567 LGM (tang2024lgm) 17.04 0.8173 0.2274 Ours 19.31 0.8238 0.1412
@@ 5.3. Ablation study
Besides the 4 input image settings, we also train our model with 6 input images and generate 3D Gaussians under the supervision of 8 images. As illustrated in Figure 5, our method achieves faithful 3DGS reconstruction using both 4 and 6 input images. Notably, qualitative results with 6 input images demonstrate greater consistency, indicating that leveraging additional images enhances reconstruction quality. Previous methods typically use a fixed number of input images (szymanowicz2024splatter; tang2024lgm). In contrast, our model can reconstruct 3D objects with a larger number of input images, allowing more synthetic images as input to enhance the overall reconstruction quality. We provide more details about the qualitative and quantitative results in the supplementary materials.
@@
We compare the training efficiency with LGM (tang2024lgm). Our experiments are conducted using our default training resources, which include 8 NVIDIA V100 (32GB) GPUs, and we train on the Objaverse (deitke2023objaverse) LVIS subset. To make a fair comparison, we set the input to 4 images and set the output Gaussians resolution to $256\times 256=65536$ for both methods. Specifically, we add one more upsampling layer for LGM to match the input resolution. As shown in Figure 6 (a), our method converges within 200 epochs. In contrast, LGM (tang2024lgm) still requires significantly more training time to reach convergence.
Additionally, we experimented with a smaller output Gaussian resolution of $128\times 128=16384$, which follows the original asymmetric design of LGM (tang2024lgm). The smaller size of predicted 3D Gaussians enables a more effective batch size, allowing our method to converge faster within around 100 epochs. Meanwhile, LGM(tang2024lgm) still requires more than 200 epochs to complete training. Both experiments indicate that our method successfully utilizes the disentangled framework to accelerate the reconstruction model regression.